A rectangle has a fixed perimeter of 350 but the length and width of the rectangle can vary. Define a function g that determines the area of the rectangle (in square inches) given the length of the rectangle in inches, l.
Answer by stanbon(75887)

A rectangle has a fixed perimeter of 350 but the length and width of the rectangle can vary. Define a function g that determines the area of the rectangle (in square inches) given the length of the rectangle in inches, l.
-----
If perimeter is 350 length + width = 350/2 = 175
Width = 175-Length
-----
Area = length*width = L(175-L)
----
Ans:: A(L) = 175L-L^2
-------
Cheers,
Stan H.
-------